16. How many years will take $25,000 to grow to $120,000 if bank offered rate is 18%. Page 1

plz answer step by step for my understanding.
Answer #1

Ans 9.48 years

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
120000 / 25000 = (1 + 18%)^n
4.8 = (1 + 18%)^n
4.8 = 1.18^n
n = 9.48 years